# TabICL: A Tabular Foundation Model for In-Context Learning on Large Data

### v1

@@ Abstract
The long-standing dominance of gradient-boosted decision trees on tabular data is currently challenged by tabular foundation models using In-Context Learning (ICL): setting the training data as context for the test data and predicting in a single forward pass without parameter updates. While the very recent TabPFNv2 foundation model (2025) excels on tables with up to 10K samples, its alternating column- and row-wise attentions make handling large training sets computationally prohibitive. So, can ICL be effectively scaled and deliver a benefit for larger tables? We introduce TabICL, a tabular foundation model for classification, pretrained on synthetic datasets with up to 60K samples and capable of handling 500K samples on affordable resources. This is enabled by a novel two-stage architecture: a column-then-row attention mechanism to build fixed-dimensional embeddings of rows, followed by a transformer for efficient ICL. Across 200 classification datasets from the TALENT benchmark, TabICL is on par with TabPFNv2 while being systematically faster (up to 10 times), and significantly outperforms all other approaches. On 56 datasets with over 10K samples, TabICL surpasses both TabPFNv2 and CatBoost, demonstrating the potential of ICL for large data. Inference code and pre-trained models are available at https://github.com/soda-inria/tabicl.
@@ 1 Introduction
While LLMs are primarily designed to model natural language, attempts to fine-tune them for tabular data have emerged (Hegselmann et al., 2023; Fang et al., 2024, and references therein). These efforts rely on table serialization, which is the process of converting table rows into text or sentences which can then be tokenized. Gardner et al. (2024) fine-tuned a Llama 3-8B on a corpus of serialized tables and demonstrated the effectiveness of this approach compared to tree-based models in the few-shot setting. However, such language models-based approaches are limited by the size of their context window to fit large serialized tables (e.g. up to 32 or 64 shots in Gardner et al. 2024). Moreover, it is unclear whether LLMs can effectively handle numerical values (Thawani et al., 2021). Finally, as evidence shows that LLMs are pretrained on many popular datasets (Bordt et al., 2024), their evaluation on tabular prediction tasks should also be conducted with care.
@@
Taking a markedly different approach, Hollmann et al. (2022) introduced TabPFN, a transformer-based tabular foundation model for classification tasks, pretrained on synthetic tabular datasets only. A key feature of TabPFN is ICL with tables, which eliminates the need for tokenization and enables efficient handling of small tables with up to 1K samples and 100 features. Very recently, the same authors introduced TabPFNv2 (2025), an improved version that significantly outperforms tree-based and neural network competitors on small-to-medium datasets with up to 10K samples and 500 features. The great promise of tabular ICL has spurred a new line of research (see Section 2.3), yet the quadratic cost of self attention is a threat to scalability of all these. TabPFNv2 uses a two-way attention mechanism alternating between column-wise and row-wise attentions, which limits its scalability for large datasets. In real-world scenarios, where industrial datasets can contain millions of samples (Rubachev et al., 2024), the high computational and memory demands of TabPFNv2 hinder its practicality.
@@
In this paper, we introduce TabICL, a scalable and efficient tabular foundation model designed for classification tasks. Pretrained on synthetic datasets with up to 60K samples, TabICL can effectively handle datasets with up to 500K samples and 500 features, significantly expanding the scalability boundaries of ICL for tables and cementing it as a foundational technique for tabular foundation models.
@@
To handle tables of arbitrary sizes with architectural changes, TabICL treats individual cells as basic elements: each column is seen as a set of cell values capturing feature-specific distribution and semantics, and each row consists of interdependent feature values providing a holistic view of each sample. TabICL employs a two-stage architecture to achieve efficient ICL for tabular data. First, it encodes rows (excluding target labels) into dense vector embeddings. Each embedding is designed to be capable of capturing the entire table information. This stage effectively collapses the column dimension to substantially reduce computational complexity and memory footprint for subsequent ICL. Second, it combines these compact yet informative embeddings with corresponding labels and then performs ICL. Therefore, the core of TabICL lies in its embedding strategy of the first stage, which is supposed to transform rows into semantically rich embeddings.
@@
Words and phrases in text often carry clear semantics, and are thus naturally associated to informative embeddings (Mikolov et al., 2013). However, tabular data lacks such an inherent structure: cell values can be ambiguous without metadata such as column names or data types. To address this challenge, TabICL adopts a well-constrained embedding strategy combining (1) distribution-aware column-wise feature embedding to capture statistical regularities within each column and (2) attention-based row-wise interaction to model dependencies across columns, thereby constructing semantically grounded representations for tabular data.
@@
In addition to the above innovations, we introduce several improvements: (1) We refine pretraining synthetic datasets of TabPFN by adding a new tree-based data-generating model to incorporate the inductive biases of tree-based models (Grinsztajn et al., 2022; den Breejen et al., 2024); (2) We adopt curriculum learning by gradually scaling the pretraining dataset size from 1K to 60K; (3) To address classification problems with over 10 classes (the pretraining limit), we use hierarchical classification (Silla & Freitas, 2011), breaking them into a hierarchical structure of subproblems with $\leq$ 10 classes. The increase in the number of tasks is largely offset by the fast ICL-powered predictions of TabICL.
To summarize our contributions: (1) We present TabICL, a novel scalable tabular foundation model for classification tasks that can accommodate any number of samples, features, and classes. In practice, TabICL handles up to 500K samples and 500 features with around 20GB of GPU memory; (2) We introduce a distribution-aware feature embedding approach that handles features with diverse properties in a unified manner, unlocking new possibilities for cross-table transferability; (3) TabICL performs tasks in a single forward pass and is orders of magnitude faster than tabular methods requiring hyper-parameter tuning while still providing better performance in most cases. TabICL is also consistently faster than TabPFNv2 (up to 10 times), with efficiency gains increasing as dataset size grows; (4) We evaluate TabICL on the TALENT benchmark (Ye et al., 2025), comprising 200 classification datasets across various domains and sizes (up to 150K samples). TabICL performs comparably to TabPFNv2 on medium-sized datasets and significantly outperforms all other methods. On the 55 large datasets with over 10K samples, TabICL surpasses both TabPFNv2 and CatBoost (Dorogush et al., 2018). These results demonstrate the potential of ICL for large data.
@@ 2 Related Work
Figure 1: An overview of the architecture of TabICL. First, column-wise embedding transforms each cell of the input table into an embedding vector using a transformer $\text{TF}_{\text{col}}$ (Section 3.2), producing $E$. Next, row-wise interaction prepends four trainable [CLS] tokens to $E$, applies rotary positional encoding, and processes $E$ row-by-row with a transformer $\text{TF}_{\text{row}}$. Concatenating the outputs of [CLS] tokens yields the final row-wise embeddings $H$. Finally, dataset-wise ICL operates on $H$ and uses a Transformer $\text{TF}_{\text{icl}}$ to predict the target labels for the test set in a single forward pass. Overall, TabICL consists of three transformers.
@@ 2.1 Foundation Models and In-Context Learning
In recent years, deep learning (DL) has been transformed by the emergence of foundation models, which are pretrained on massive, diverse datasets and serve as versatile backbones for downstream tasks. These transformer-based models enable In-Context Learning (ICL): performing tasks by analyzing prompts containing input-output pairs, without explicit training or parameter updates. ICL operates as a form of on-the-fly reasoning. The mechanism underlying ICL remains elusive, with prevailing explanations framing it as implicit Bayesian inference (Xie et al., 2022; Müller et al., 2024), gradient descent optimization (von Oswald et al., 2023), and algorithmic learning (Garg et al., 2023).
@@ 2.2 Tabular Deep Learning Models
Gradient-boosted decision trees (GBDTs), such as CatBoost and XGBoost (Chen & Guestrin, 2016), have long dominated the tabular domain. However, there are growing efforts to develop tabular DL models. Recent studies indicate a narrowing performance gap between GBDTs and tabular DL models (Ye et al., 2024; Gorishniy et al., 2024).
As tabular DL improves, cross-table transferability emerges as an important topic. Notable efforts in this direction include XTab (Zhu et al., 2023) and CARTE (Kim et al., 2024), which incorporate transferable components that are typically shareable backbone networks and dataset-specific components that require fine-tuning for each new task. The advent of tabular foundation models can bring new possibilities to cross-table learning, paving the way for large-scale pretraining and transfer learning across tables.
2.3 TabPFN and its Offsprings
@@
Other approaches try to improve the quality of pre-training data, such as TabForestPFN (den Breejen et al., 2024) incorporating tree-based synthetic datasets and TabDPT (Ma et al., 2024a) curating and using real-world datasets.
Very recently (January 2025), TabPFNv2 was released and largely improved TabPFN in terms of both prediction performance and scalability. Our contributed model, TabICL, achieves comparable performance to TabPFNv2 while being more scalable and computationally efficient.
@@ 3 The TabICL Architecture
We consider a tabular classification task with an input space $\mathcal{X}\in\mathbb{R}^{m}$ and a target space $\mathcal{Y}\in[1,\cdots,C]$. Given a training dataset of input-output pairs $\mathcal{D}_{\text{train}}=\{(x_{\text{train}}^{(i)},y_{\text{train}}^{(i)})\}_{i=1}^{n_{\text{tr}}}$ and test samples $X_{\text{test}}=\{x_{\text{test}}^{(i)}\}_{i=1}^{n_{\text{te}}}$, our goal is to predict the class probabilities $p(\cdot|x_{\text{test}},\mathcal{D}_{\text{train}})$.
@@
(b) Column-wise embedding
@@ 3.2 Distribution-aware Column-wise Embedding
The column-wise embedding (or feature embedding) maps each scalar cell in a column $c_{j}\in\mathbb{R}^{n}$ into a $d$-dimensional embedding. Unlike typical approaches that assign a separate linear layer to each column (Gorishniy et al., 2022), we embed all columns through a shareable Set Transformer $\text{TF}_{\text{col}}$, as illustrated in Figure 2(b) and formulated as follows:
@@
Note that each cell in a column is assigned its own weight and bias. Essentially, feature embedding can be framed as a set-input problem, where $\text{TF}_{\text{col}}$ serves as a hypernetwork taking as input a permutation-invariant set of cell values and generating distribution-aware weights and biases.
@@
where Lin represents a linear layer, MAB denotes multi-head attention block, and IASB stands for induced self-attention block (Figure 2(a)), introduced by (Lee et al., 2019). A column $c$ (omitting the subscript $j$ for simplicity) is projected into a $d$-dimensional space ($d=128$) via a linear layer, processed by ISAB consisting of $\text{MAB}_{1}$ and $\text{MAB}_{2}$, and passed through linear layers to generate $W$ and $B$.
@@
To get insights into the information captured in the learned feature embeddings, we visualize $M\in\mathbb{R}^{k\times d}$ (Equation 4) produced by the final ISAB block of $\text{TF}_{\text{col}}$. We summarize $M$ by summing along the first dimension (i.e., aggregating the induced representations of all inducing vectors) to obtain a single vector per column, and then apply Principal Component Analysis. Figure 3 reveals that columns with similar skewness (resp. kurtosis) tend to cluster together, i.e., non-symmetric distributions are differentiated from symmetrical ones (Figure 3 left), and heavy-tailed distributions are differentiated from light-tailed ones. This suggests that $\text{TF}_{\text{col}}$ encodes distributional properties in a structured manner, and cells in a column are probably embedded to reflect their statistical role (e.g., min, max, mean, mode). Therefore, the learned feature embeddings should distinguish features based on their unique distributional properties, effectively serving as feature identifiers. This contrasts with methods that rely on semantic column names (Kim et al., 2024) or learn feature identifier vectors (Kossen et al., 2021).
@@ 3.3 Context-aware Row-wise Interaction
A defining characteristic of tabular data is that columns do not have a natural ordering. Ideally, tabular methods should be invariant to permutations of columns. Unlike TabPFN (v1), TabICL naturally incorporates this invariance through the row-wise attention. However, we experimentally observed that $\text{TF}_{\text{row}}$ can suffer from a representation collapse issue. As mentioned earlier, features are identified by their distributional properties after column-wise embedding. Consequently, features originating from similar distributions thus become less distinguishable. In the extreme case where all features are drawn from the same distribution, $\text{TF}_{\text{row}}$ cannot differentiate a sample from any of its column-permuted versions, leading to nearly identical representations for originally distinct samples, i.e., representation collapse. This phenomenon is exemplified by the balance scale dataset (Siegler, 1976), as shown in Figure 4. In this dataset, all features follow the same discrete distribution and can take only 5 values, making collapse highly probable. After processing with $\text{TF}_{\text{row}}$, we observe that many samples collapse to the same representation (rightmost plot), despite being originally distinct (leftmost plot).
To break the symmetry between identically distributed features, we incorporate rotary positional embedding (RoPE, Su et al., 2023) in $\text{TF}_{\text{row}}$. RoPE is widely adopted in recent LLMs (Dubey et al., 2024) and directly encodes relative positional information into the attention mechanism by rotating the query and key vectors. The rotation angle is determined by the position $p$ in the sequence and the dimension index $i$, defined as $\theta_{i}=p/(\text{base}^{2i/d})$, where $d$ is the embedding dimension and base is the frequency scaling factor. More details can be found in Appendix C.
@@
While the use of RoPE breaks permutational invariance, it can mitigate the representation collapse. For example, on the balance scale dataset, RoPE preserves distinct representations for different samples (Figure 4 middle). Following Xiong et al. (2023), we set a large scaling factor of 100,000 for RoPE to enhance generalization to a larger number of features than those seen during training (up to 100). To approximately restore permutation invariance, TabICL adopts the same strategy as other TabPFN-like models by ensembling predictions over multiple column permutations.

### v2

@@ Abstract
The long-standing dominance of gradient-boosted decision trees on tabular data is currently challenged by tabular foundation models using In-Context Learning (ICL): setting the training data as context for the test data and predicting in a single forward pass without parameter updates. While TabPFNv2 foundation model excels on tables with up to 10K samples, its alternating column- and row-wise attentions make handling large training sets computationally prohibitive. So, can ICL be effectively scaled and deliver a benefit for larger tables? We introduce TabICL, a tabular foundation model for classification, pretrained on synthetic datasets with up to 60K samples and capable of handling 500K samples on affordable resources. This is enabled by a novel two-stage architecture: a column-then-row attention mechanism to build fixed-dimensional embeddings of rows, followed by a transformer for efficient ICL. Across 200 classification datasets from the TALENT benchmark, TabICL is on par with TabPFNv2 while being systematically faster (up to 10 times), and significantly outperforms all other approaches. On 53 datasets with over 10K samples, TabICL surpasses both TabPFNv2 and CatBoost, demonstrating the potential of ICL for large data. Pretraining code, inference code, and pre-trained models are available at https://github.com/soda-inria/tabicl.
@@ 1 Introduction
While LLMs are primarily designed to model natural language, attempts to fine-tune them for tabular data have emerged (Hegselmann et al., 2023; Fang et al., 2024, and references therein). These efforts rely on table serialization, which is the process of converting table rows into text or sentences which can then be tokenized. Gardner et al. (2024) fine-tuned Llama 3-8B on a corpus of serialized tables and demonstrated the effectiveness of this approach compared to tree-based models in the few-shot setting. However, such language models-based approaches are limited by the size of their context window to fit large serialized tables (e.g. up to 32 or 64 shots in Gardner et al. 2024). Moreover, it is unclear whether LLMs can effectively handle numerical values (Thawani et al., 2021). Finally, as evidence shows that LLMs are pretrained on many popular datasets (Bordt et al., 2024), their evaluation on tabular prediction tasks should also be conducted with care.
@@
Taking a markedly different approach, Hollmann et al. (2022) introduced TabPFN, a transformer-based tabular foundation model for classification tasks, pretrained on synthetic tabular datasets only. A key feature of TabPFN is ICL with tables, which eliminates the need for tokenization and enables efficient handling of small tables with up to 1K samples and 100 features. The same authors then introduced TabPFNv2 (Hollmann et al., 2025), an improved version that significantly outperforms tree-based and neural network competitors on small-to-medium datasets with up to 10K samples and 500 features. The great promise of tabular ICL has spurred a new line of research (see Section 2.3), yet the quadratic cost of self-attention is a threat to the scalability of all of these. TabPFNv2 uses a two-way attention mechanism alternating between column-wise and row-wise attentions, which limits its scalability for large datasets. In real-world scenarios, where industrial datasets can contain millions of samples (Rubachev et al., 2024), the high computational and memory demands of TabPFNv2 hinder its practicality.
@@
In this paper, we introduce TabICL, a scalable and efficient tabular foundation model based on PFNs (Müller et al., 2024) and designed for classification tasks. Pretrained on synthetic datasets up to 60K samples, TabICL can effectively handle datasets up to 500K samples and 500 features, significantly expanding the scalability of ICL for tables.
@@
To accommodate tables of arbitrary sizes, TabICL treats individual cells as fundamental units: each column is regarded as a set of cell values that capture feature-specific distributions and semantics, while each row comprises interdependent feature values. TabICL employs a two-stage architecture to achieve efficient ICL for tabular data. First, it encodes rows (excluding target labels) into dense vector embeddings. Each embedding is designed to capture the entire table information. This stage effectively collapses the column dimension to substantially reduce computational complexity and memory footprint for subsequent ICL. Second, it combines these compact yet informative embeddings with corresponding labels and then performs ICL. Therefore, the core of TabICL lies in its embedding strategy of the first stage, which is supposed to transform rows into semantically rich embeddings.
@@
In natural language, words often carry clear semantics and are thus naturally associated to informative embeddings (Mikolov et al., 2013). However, tabular data lacks such an inherent structure: cell values can be ambiguous without metadata such as column names or data types. To address this challenge, TabICL adopts a well-constrained embedding strategy combining (1) distribution-aware column-wise embedding to capture statistical regularities within each column and (2) attention-based row-wise interaction to model dependencies across columns, thereby constructing semantically grounded representations for tabular data.
@@
In addition to the above innovations, we introduce further improvements: (1) We refine TabPFN’s pretraining synthetic datasets by adding a new tree-based data-generating model to incorporate the inductive biases of tree-based models (Grinsztajn et al., 2022; den Breejen et al., 2024; Grinsztajn, 2024); (2) We adopt curriculum learning by scaling the pretraining dataset size from 1K to 60K; (3) To address classification problems with over 10 classes (the pretraining limit), we use hierarchical classification (Silla & Freitas, 2011), breaking them into a hierarchical structure of subproblems with $\leq$ 10 classes. The increase in the number of tasks is largely offset by the fast predictions of TabICL.
To summarize our contributions: (1) We present TabICL, a novel scalable tabular foundation model for classification tasks that can accommodate any number of samples, features, and classes. In practice, TabICL handles up to 500K samples and 500 features with around 20GB of GPU memory; (2) We introduce a distribution-aware feature embedding approach that handles features with diverse properties in a unified manner, unlocking new possibilities for cross-table transferability; (3) TabICL performs tasks in a single forward pass and is orders of magnitude faster than tabular methods requiring hyperparameter tuning while still providing better performance in most cases. TabICL is also consistently faster than TabPFNv2 (up to 10 times), with efficiency gains increasing as dataset size grows; (4) We evaluate TabICL on the TALENT benchmark (Ye et al., 2025), comprising 200 classification datasets across various domains and sizes (up to 150K samples). TabICL performs comparably to TabPFNv2 on medium-sized datasets and significantly outperforms all other methods. On the 53 large datasets with over 10K samples, TabICL surpasses both TabPFNv2 and CatBoost (Dorogush et al., 2018). These results demonstrate the potential of ICL for large data.
@@ 2 Related Work
Figure 1: Architecture Overview of TabICL. First, column-wise embedding transforms each cell of the input table into an embedding vector using a transformer $\text{TF}_{\text{col}}$ (Section 3.2), producing $E$. Next, row-wise interaction prepends four trainable [CLS] tokens to $E$, applies rotary positional encoding, and processes $E$ row-by-row with a transformer $\text{TF}_{\text{row}}$. Concatenating the outputs of [CLS] tokens yields the final row-wise embeddings $H$. Finally, dataset-wise ICL operates on $H$ and uses a Transformer $\text{TF}_{\text{icl}}$ to predict the target labels for the test set in a single forward pass. Overall, TabICL consists of three transformers.
@@ 2.1 Foundation Models and In-Context Learning
In recent years, deep learning has been transformed by the emergence of foundation models, which are pretrained on massive, diverse datasets and serve as versatile backbones for downstream tasks. These transformer-based models enable In-Context Learning (ICL): performing tasks by analyzing prompts containing input-output pairs, without explicit training or parameter updates. ICL operates as a form of on-the-fly reasoning. The mechanism underlying ICL remains elusive, with prevailing explanations framing it as implicit Bayesian inference (Xie et al., 2022; Müller et al., 2024), gradient descent optimization (von Oswald et al., 2023), and algorithmic learning (Garg et al., 2023).
@@ 2.2 Tabular Deep Learning Models
Gradient-boosted decision trees (GBDTs), such as CatBoost and XGBoost (Chen & Guestrin, 2016), have long dominated the tabular domain. However, growing efforts are focused on developing deep learning models for tabular data. Recent studies indicate a narrowing performance gap between GBDTs and tabular deep learning models (Ye et al., 2024; Gorishniy et al., 2024).
As tabular deep learning improves, cross-table transferability emerges as an important topic. Notable efforts in this direction include XTab (Zhu et al., 2023) and CARTE (Kim et al., 2024), which incorporate transferable components that are typically shareable backbone networks and dataset-specific components that require fine-tuning for each new task. The advent of tabular foundation models can bring new possibilities to cross-table learning, paving the way for large-scale pretraining and transfer learning across tables.
2.3 Tabular Foundation Models
@@
Several TabPFN variants try to improve the quality of pre-training data, such as TabForestPFN (den Breejen et al., 2024) incorporating tree-based synthetic datasets and TabDPT (Ma et al., 2024a) using real-world datasets.
TabPFNv2 largely improves TabPFN in terms of both prediction performance and scalability. Our new model, TabICL, achieves comparable performance to TabPFNv2 while being more scalable and computationally efficient. Appendix A provides a systematic comparison between TabPFNv2 and TabICL across architecture, pretraining, and scalability.
@@ 3 The TabICL Architecture
We consider a tabular classification task with an input space $\mathcal{X}\in\mathbb{R}^{m}$ and a target space $\mathcal{Y}\in[1,\cdots,C]$. Given a training dataset of input-output pairs $\mathcal{D}_{\text{train}}=\{(x_{\text{train}}^{(i)},y_{\text{train}}^{(i)})\}_{i=1}^{n_{\text{tr}}}$ and test sample $x_{\text{test}}$, our goal is to predict the class probabilities $p(\cdot|x_{\text{test}},\mathcal{D}_{\text{train}})$.
@@
(b) Column-wise inter-sample embedding
@@ 3.2 Distribution-aware Column-wise Embedding
The column-wise embedding (or feature embedding) maps each scalar cell in a column $c_{j}\in\mathbb{R}^{n}$ into a $d$-dimensional embedding. Unlike typical approaches that assign a separate embedding module to each column (Gorishniy et al., 2022), we embed all columns through a shareable Set Transformer $\text{TF}_{\text{col}}$, as illustrated in Figure 2(b) and formulated as follows:
@@
Note that each cell in a column is assigned its own weight and bias. Essentially, feature embedding can be framed as a set-input problem, where $\text{TF}_{\text{col}}$ serves as a hypernetwork taking as input a permutation-invariant set of cell values and generating distribution-aware weights and biases $W$ and $B$.
@@
where Lin represents a linear layer, MAB denotes a multi-head attention block, and IASB stands for induced self-attention block (Figure 2(a)), introduced by (Lee et al., 2019). A column $c$ (omitting the subscript $j$ for simplicity) is projected into a $d$-dimensional space ($d=128$) via a linear layer, processed by ISAB consisting of $\text{MAB}_{1}$ and $\text{MAB}_{2}$, and passed through linear layers to generate $W$ and $B$.
@@
To get insights into the information captured in the learned feature embeddings, we visualize $M\in\mathbb{R}^{k\times d}$ (Equation 4) the output of the final ISAB block of $\text{TF}_{\text{col}}$. We summarize $M$ by summing along the first dimension (i.e., aggregating the induced representations of all inducing vectors) to obtain a single vector per column, and then apply Principal Component Analysis. Figure 3 reveals that columns with similar skewness (resp. kurtosis) tend to cluster together, i.e., non-symmetric distributions are differentiated from symmetrical ones (Figure 3 left), and heavy-tailed distributions are differentiated from light-tailed ones. This suggests that $\text{TF}_{\text{col}}$ encodes distributional properties in a structured manner, and cells in a column are probably embedded to reflect their statistical role (e.g., min, max, mean, mode). Therefore, the learned feature embeddings should distinguish features based on their unique distributional properties, effectively serving as feature identifiers. This contrasts with methods that rely on semantic column names (Kim et al., 2024) or learn feature identifier vectors (Kossen et al., 2021).
@@ 3.3 Context-aware Row-wise Interaction
We experimentally observed that $\text{TF}_{\text{row}}$ can suffer from a representation collapse issue. As mentioned earlier, features are identified by their distributional properties after column-wise embedding. Consequently, features originating from similar distributions thus become less distinguishable. In the extreme case where all features are drawn from the same distribution, $\text{TF}_{\text{row}}$ using the vanilla permutation-invariant self-attention cannot differentiate a sample from any of its column-permuted versions, leading to nearly identical representations for originally distinct samples, i.e., representation collapse. This phenomenon is exemplified by the balance scale dataset (Siegler, 1976), as shown in Figure 4. In this dataset, all features follow the same discrete distribution with only 5 possible values, making collapse highly probable. After processing with $\text{TF}_{\text{row}}$, we observe that many samples collapse to the same representation (rightmost plot), despite being originally distinct (leftmost plot).
A similar issue of representation collapse was reported in TabPFNv2. To mitigate this, Hollmann et al. (2025) introduced random feature identifier vectors for each feature and encoded groups of features collectively rather than individually. In contrast, we employ a different strategy by incorporating rotary positional embedding (RoPE, Su et al., 2023) into $\text{TF}_{\text{row}}$ to break the symmetry between identically distributed features. RoPE is widely adopted in recent LLMs (Dubey et al., 2024) and directly encodes relative positional information into the attention mechanism by rotating the query and key vectors. The rotation angle is determined by the position $p$ in the sequence and the dimension index $i$, defined as $\theta_{i}=p/(\text{base}^{2i/d})$, where $d$ is the embedding dimension and base is the frequency scaling factor. More details can be found in Appendix D.
@@
While the use of RoPE breaks permutation invariance with respect to column order, it can effectively mitigate the representation collapse. For instance, on the balance scale dataset, RoPE preserves distinct representations across different samples (Figure 4 middle). Following Xiong et al. (2023), we use a large scaling factor of 100,000 for RoPE to enhance generalization to a greater number of features than seen during training (up to 100). To approximately restore permutation invariance, TabICL adopts the same strategy as other TabPFN-like models by ensembling predictions over multiple column permutations.